# STeP: A General and Scalable Framework for Solving Video Inverse Problems with Spatiotemporal Diffusion Priors

### v1

@@ Abstract
We study how to solve general Bayesian inverse problems involving videos using diffusion model priors. While it is desirable to use a video diffusion prior to effectively capture complex temporal relationships, due to the computational and data requirements of training such a model, prior work has instead relied on image diffusion priors on single frames combined with heuristics to enforce temporal consistency. However, these approaches struggle with faithfully recovering the underlying temporal relationships, particularly for tasks with high temporal uncertainty. In this paper, we demonstrate the feasibility of practical and accessible spatiotemporal diffusion priors by fine-tuning latent video diffusion models from pretrained image diffusion models using limited videos in specific domains. Leveraging this plug-and-play spatiotemporal diffusion prior, we introduce a general and scalable framework for solving video inverse problems. We then apply our framework to two challenging scientific video inverse problems—black hole imaging and dynamic MRI. Our framework enables the generation of diverse, high-fidelity video reconstructions that not only fit observations but also recover multi-modal solutions. By incorporating a spatiotemporal diffusion prior, we significantly improve our ability to capture complex temporal relationships in the data while also enhancing spatial fidelity. Our code is available at the GitHub repository STeP.
Figure 1: An overview of our proposed framework with Spatiotemporal Diffusion Priors (STeP) for video inverse problems. Left: We consider solving video inverse problems in scientific domains where a relatively large image dataset is available for training a prior but the amount of video data is limited. Middle: We propose a scalable and data-efficient spatiotemporal diffusion prior that directly models the video distribution using collections of both images and videos (samples generated from the prior are shown in the light purple box). We combine the prior and knowledge of the inverse problem in a state-of-the-art plug-and-play (PnP) diffusion solver [81]. Right: The resulting algorithm can recover multi-modal posterior distributions for difficult ill-posed inverse problems. Here we demonstrate the approach on the black hole video reconstruction problem, a highly nonlinear inverse problem with extremely sparse measurements leading to a multi-modal posterior. Specifically, we generated 100 posterior video samples and observe that there are three modes with equal data fidelity but with significantly different spatiotemporal structures (as shown by the frames and average optical flow from [59] in the last column). One of the recovered modes matches the ground truth in both spatial and temporal structure. This example shows the capability of our framework to generate diverse and high-fidelity video reconstructions for challenging scientific problems.
**footnotetext: These authors contributed equally to this work
@@ 1 Introduction
Using diffusion models as priors for solving Bayesian inverse problems has emerged as a powerful approach, demonstrating remarkable effectiveness in imposing image statistics learned from training data to guide recovered solutions. Plug-and-play (PnP) inversion methods that make use of diffusion priors have been successfully applied to diverse applications, including image restoration [67, 54, 34, 15, 48, 40, 87], medical imaging [30, 55, 13, 29, 18, 14, 81], and physics-based inverse problems [84, 1, 24, 56, 57, 73]. As a PnP prior, a diffusion model can be applied to various problems without retraining the model, making it flexible and easy to use. The success of these methods relies on two key factors: (1) access to a well-trained diffusion model, learned from a large set of unlabeled source data, and (2) a robust PnP framework capable of handling inverse problems with different underlying challenges [73, 81, 85, 84, 12].
Most prior work has focused on solving inverse problems for images; however, many critical inverse problems inherently involve temporal information, necessitating a general framework for solving video inverse problems [7, 65, 44, 79, 42]. Because training a video diffusion model is commonly believed to be too computationally challenging and data hungry, existing approaches to video inverse problems rely on image diffusion priors [17, 37, 36, 77], which process each frame independently together with various heuristics based on correlated noise or optical flow information to enforce temporal consistency (see Fig. 2 for a schematic illustration). However, these methods struggle to faithfully recover complex temporal relationships when observations become sparse and ill-posed, which is common in scientific inverse problems [65, 44].
@@
In this paper, we propose a general and scalable approach for addressing video inverse problems, STeP, by integrating a SpatioTemporal video diffusion Prior into a PnP inversion method. To do so, we first demonstrate the feasibility of training a video diffusion prior for solving inverse problems using a limited amount of video data. Instead of training a video diffusion model from scratch, inspired by [66], we start from an image diffusion model and fine-tune the temporal modules, transforming it into a spatiotemporal video diffusion model using only a few hundred to a few thousand videos. This approach enables video diffusion in a data-efficient manner, drastically reducing training cost and making it feasible to obtain a video diffusion model from an image diffusion model within just a few hours on a single A100 GPU. After obtaining a well-trained spatiotemporal diffusion prior, we integrate it with a state-of-the-art PnP inversion method, namely the Decoupled Annealing Posterior Sampling (DAPS) [81] framework. STeP inherits the ability of DAPS to handle general inverse problems (with nonlinear forward models) and does not require additional temporal heuristics for solving video inverse problems.
@@
We demonstrate the effectiveness of STeP on two challenging scientific video inverse problems: black hole video reconstruction (previewed in Fig. 1) and dynamic magnetic resonance imaging (MRI). Our experiments show that a fine-tuned spatiotemporal diffusion prior can be seamlessly integrated with the existing PnP diffusion solver, enabling efficient posterior sampling. As Fig. 1 illustrates, STeP not only achieves state-of-the-art results with improved temporal and spatial consistency but also effectively captures the multi-modal nature of highly ill-posed problems, recovering diverse plausible solutions from the posterior distribution. Notably, it achieves substantial improvements in temporal consistency, with a 6.50dB and 2.69dB increase in d-PSNR (average PSNR of difference images between all consecutive frames of a video) for black hole imaging and dynamic MRI, respectively—where d-PSNR quantifies temporal coherence. STeP also outperforms baselines in terms of spatial consistency by 1.69dB and 1.15dB in average frame-wise PSNR for black hole imaging and dynamic MRI, respectively.
@@
Figure 2: A schematic comparison between prior works (top) and our STeP framework (bottom) for video inverse problems. The bold texts highlight the key differences between them. While prior works only use an image diffusion model and enforce temporal consistency with various heuristics, we directly learn a spatiotemporal diffusion prior. By leveraging a spatiotemporal prior, we improve both the temporal consistency and per-frame spatial consistency of the generated videos within a general and scalable PnP diffusion framework.
@@ 2 Background
2.1 Video latent diffusion models
Diffusion models [51, 53, 27, 54, 31] generate data by reversing a predefined noising process. Starting from the data distribution $p({\mathbf{x}}_{0})$, noisy data distributions $p({\mathbf{x}}_{t};\sigma_{t})$ are created by adding Gaussian noise with standard deviation $\sigma_{t}$, where $\sigma_{t}$ is a predefined noise schedule. To sample from the diffusion model, one requires the time-dependent score function $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})$ [51, 54, 31], which can be approximated by training a network ${\mathbf{s}}_{\bm{\theta}}({\mathbf{x}}_{t},\sigma_{t})$ using denoising score matching [64] with either a UNet [27, 32] or a transformer [43, 76] architecture.
While training diffusion models on the original high-dimensional data space may suffer from high computational cost, latent diffusion models (LDM) [46] instead generate an efficient, low-dimensional latent representation ${\mathbf{z}}_{0}$ of data ${\mathbf{x}}_{0}$ with a pretrained perceptual compression encoder $\mathcal{E}$ and decoder $\mathcal{D}$, which satisfy ${\mathbf{z}}_{0}=\mathcal{E}({\mathbf{x}}_{0})$ and $\mathcal{D}({\mathbf{z}}_{0})\approx{\mathbf{x}}_{0}$. The compression models $\mathcal{E}$ and $\mathcal{D}$ can be trained with VAE variants [35, 45, 8] with KL divergence regularization or VQGAN variants [26, 21, 63] with quantization regularization.
Video latent diffusion models are commonly believed to be hard to train due to computational cost in 3D modules in architecture and the requirement of a large video dataset [6, 86, 83, 28, 5]. Many recent methods tend to solve video modeling by training or fine-tuning from a pretrained image diffusion model with a video dataset.
2.2 Inverse problems with diffusion priors
Various methods have been proposed to solve Bayesian inversion using a pretrained diffusion model as a prior. Guidance-based methods [15, 47, 34, 50] approximate the intractable noisy likelihood score term $\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{y}}|{\mathbf{x}}_{t})$ while solving the reverse diffusion process. Variable splitting methods [16, 39, 49, 74, 73, 81] decompose inference into two alternating steps: one for enforcing the prior and another for incorporating the likelihood. Variational Bayes approaches [41, 24, 23] introduce a parameterized distribution to directly learn the posterior $p({\mathbf{x}}_{0}|{\mathbf{y}})$ with a diffusion prior. Sequential Monte Carlo (SMC) methods [71, 60, 10, 19] integrate diffusion sampling with SMC techniques to provide asymptotic convergence guarantees.
These methods rely on different requirements for the inverse problem and the prior. Some are specifically designed for linear problems [19, 37, 47], while others can handle nonlinear ones [73, 2]. Certain approaches require specialized designs for latent diffusion models [47, 49], whereas others can be naturally applicable [81, 39]. These requirements influence their generalizability and scalability for scientific inverse problems in different domains.
2.3 Video Inverse Problems (VIPs)
Recently, several works have extended diffusion-based approaches to video inverse problems (VIPs) [17, 88, 36, 37]. Some methods rely on image diffusion models with manually designed strategies, such as batch-consistent sampling (BCS) [37, 36] or using optical flow estimated from observations to warp the noise [17, 77]. However, we show that BCS has limited ability to faithfully recover the underlying temporal relationships for scientific inverse problems, in which the types of measurements considered often highly corrupt the temporal information. By implicitly imposing a static temporal prior via correlated noise, BCS struggles to recover complex temporal dynamics from such corrupted measurements. Previous work also highlighted a key limitation of optical flow-based methods [17, 77] being the reliance on accurate estimation of the optical flow from measurement data. When such measurements are sparse and contain limited temporal information, these methods struggle to impose an accurate temporal prior, leading to suboptimal reconstruction. Another line of work fine-tunes image diffusion models on domain-specific datasets to improve temporal consistency [88].
Compared to image diffusion models, data-driven video diffusion models offer a more general spatiotemporal prior for solving VIPs. In this work, we show the feasibility of training video diffusion models in scientific domains and their effectiveness in tackling video inverse problems.
@@ 3 Method
In this section, we introduce our framework, STeP, for solving video inverse problems with SpatioTemporal diffusion Priors. A schematic of our framework is provided in Fig. 2. We start by introducing our problem formulation in Sec. 3.1. We then propose a scalable and data-efficient way of training spatiotemporal diffusion priors in Sec. 3.2. We finally show in Sec. 3.3 that once such a prior is trained, it can be used to solve general video inverse problems.
Notations.
We adopt the following notations throughout the rest of the paper to avoid confusion. We use the variable ${\mathbf{x}}$ to denote objects in the image/video space and variable ${\mathbf{z}}$ to denote latent codes in the latent space. The variable ${\mathbf{y}}$ is always used for the measurements. Subscript $(\cdot)_{t}$ is the time index in the context of the diffusion process, where $t=0$ indicates the clean image. Superscript $(\cdot)^{[j]}$ is the index for the $j$-th frame in a video.
3.1 Basic formulation
We consider general video inverse problems (VIPs) of recovering an underlying target ${\mathbf{x}}_{0}$ from the measurements
${\mathbf{y}}=\mathcal{A}({\mathbf{x}}_{0})+{\mathbf{n}}$ (1)
where $\mathcal{A}(\cdot)$ is the forward model and ${\mathbf{n}}$ is the measurement noise. Importantly, ${\mathbf{x}}_{0}$ evolves over time, and substantial spatiotemporal information may be lost in the measurement process due to the ill-posed nature of $\mathcal{A}(\cdot)$, making it necessary to impose a prior on both the spatial and temporal dimensions of ${\mathbf{x}}_{0}$ for meaningful recovery. Our goal is to draw samples from the posterior distribution $p({\mathbf{x}}_{0}|{\mathbf{y}})\,\propto\,p({\mathbf{y}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})$. While the likelihood $p({\mathbf{y}}|{\mathbf{x}}_{0})$ can be derived from Eq. 1, it is often challenging to characterize the prior distribution $p({\mathbf{x}}_{0})$ for videos because of their high dimensionality and potentially limited number of samples for training.
@@ 3.1 Basic formulation
To overcome the challenge of high dimensionality, we propose to impose a spatiotemporal prior in latent space. Assuming that the set of likely ${\mathbf{x}}_{0}$’s is in the range of a decoder $\mathcal{D}$, we have that $\exists\,{\mathbf{z}}_{0}$ s.t. ${\mathbf{x}}_{0}=\mathcal{D}({\mathbf{z}}_{0})$ and can thus rewrite Eq. 1 as:
@@
It follows that the posterior $p({\mathbf{x}}_{0}|{\mathbf{y}})$ is the pushforward of the latent posterior $p({\mathbf{z}}_{0}|{\mathbf{y}})$ by $\mathcal{D}$, so it suffices to first generate latent samples from $p({\mathbf{z}}_{0}|{\mathbf{y}})$ and then decode them by $\mathcal{D}$.
3.2 Spatiotemporal diffusion prior
In order to meet the challenges of real-world VIPs, we aim for spatiotemporal diffusion priors with the following three properties:
  • (P1) It should be able to model distributions of high-resolution multi-frame videos and be reasonably efficient so that repeatedly calling it in a downstream solver would be computationally tractable.
  • (P2) It should directly learn temporal information from data instead of relying on heuristics so that it can capture sophisticated temporal dynamics and relationships.
  • (P3) It should be able to learn spatial information from both videos and images, given that static images are usually much more abundant for training than videos.
Our design of spatiotemporal diffusion priors aligns with each of these three properties, as discussed below.

### v2

# STeP: A Framework for Solving Scientific Video Inverse Problems with Spatiotemporal Diffusion Priors

@@ Abstract
Reconstructing spatially and temporally coherent videos from time-varying measurements is a fundamental challenge in many scientific domains. A major difficulty arises from the sparsity of measurements, which hinders accurate recovery of temporal dynamics. Existing image diffusion-based methods rely on extracting temporal consistency directly from measurements, limiting their effectiveness on scientific tasks with high spatiotemporal uncertainty. We address this difficulty by proposing a plug-and-play framework that incorporates a learned spatiotemporal diffusion prior. Due to its plug-and-play nature, our framework can be flexibly applied to different video inverse problems without the need for task-specific design and temporal heuristics. We further demonstrate that a spatiotemporal diffusion model can be trained efficiently with limited video data. We validate our approach on two challenging scientific video reconstruction tasks—black hole video reconstruction and dynamic MRI. While baseline methods struggle to provide temporally coherent reconstructions, our approach achieves significantly improved recovery of the spatiotemporal structure of the underlying ground truth videos.
@@ 1 Introduction
Reconstructing high-quality videos from time-varying measurements is a core challenge in many scientific domains, such as black hole video reconstruction (akiyama2022firstIII, ) and dynamic magnetic resonance imaging (MRI) (gamper2008compressed, ), where recovered spatiotemporal features influence downstream scientific analysis or medical interpretation. These problems are inherently difficult due to the high dimensionality of the underlying video and the severe loss of both spatial and temporal information during acquisition.
Diffusion models (DMs) have become the state-of-the-art for modeling image and video distributions, offering powerful priors for solving inverse problems zheng2024ensemblekalmandiffusionguidance . However, most existing diffusion-based approaches for video inverse problems are motivated by either video editing tasks or restoration problems in photography (daras2024warped, ; kwon2025solving, ; kwon2024visionxlhighdefinitionvideo, ; yeh2024diffir2vrzerozeroshotvideorestoration, ), where the spatiotemporal structure is either given or mostly preserved in the measurements. These methods leverage image diffusion models (IDMs) to process each video frame independently while enforcing temporal consistency via correlated noise or optical flow. We find that these methods struggle on scientific problems where there is limited spatial structure in the measurements to capture complex dynamics.
Plug-and-play diffusion prior (PnPDP) methods offer a general strategy for solving inverse problems by leveraging pre-trained diffusion models as priors. They have shown strong performance across a range of applications, including natural image restoration (wang2022zero, ; song2021scorebased, ; kawar2022denoising, ; chung2023diffusion, ; saharia2023Image, ; lugmayr2022repaint, ; zhu2023denoising, ), medical imaging (jalal2021robust, ; song2022solving, ; chung2022score, ; hung2023med, ; dorjsembe2024conditional, ; chung2022mr, ; zhang2024improving, ), and physics-based inverse problems (zheng2024ensemblekalmandiffusionguidance, ; akiyama2019first, ; feng2023score, ; sun2020deep, ; sun2023provable, ; wu2024principled, ). Despite being compatible with videos in principle, existing PnPDP methods are predominantly investigated for static images. It remains largely unexplored how to effectively obtain a well-trained video diffusion prior and extend the PnPDP framework to ill-posed video inverse problems.
@@
In this paper, we introduce STeP, a framework for solving scientific video inverse problems by integrating a SpatioTemporal video diffusion Prior into a PnPDP method. STeP enforces learned priors over both spatial and temporal dimensions and does not rely on heuristics to ensure temporal consistency, making it well-suited for tasks with high spatiotemporal uncertainty. To obtain a spatiotemporal prior efficiently, we are inspired by (wang2023modelscopetexttovideotechnicalreport, ) to first train an image latent diffusion model with a 2D UNet and then turn it into a spatiotemporal video diffusion model by adding a zero-initialized temporal module to each 2D convolution module in the UNet. This enables us to fine-tune a spatiotemporal diffusion prior from a pre-trained IDM in a data- and time-efficient manner, using only hundreds to thousands of videos and a few hours of training on a single A100 GPU. We then integrate the resulting prior with the Decoupled Annealing Posterior Sampling (DAPS) (zhang2024improving, ), a state-of-the-art PnPDP method. STeP inherits the ability of DAPS to handle general inverse problems with nonlinear forward models and does not require heuristics to enforce temporal consistency.
@@
We demonstrate the effectiveness of STeP on two challenging scientific video inverse problems: black hole video reconstruction (previewed in Figure 1) and dynamic MRI, where the underlying targets exhibit significantly different spatiotemporal characteristics. As Figure 1 illustrates, STeP not only achieves state-of-the-art results with improved spatial and temporal consistency but also effectively captures the multi-modal nature of highly ill-posed problems, recovering diverse plausible solutions from the posterior distribution. Notably, our approach achieves substantial improvements in both spatial and temporal consistency, significantly outperforming baselines across most evaluation metrics. For example, our method demonstrates a 1.57dB and 2.38dB improvement on PSNR on black hole video reconstruction and dynamic MRI, respectively. More importantly, for the challenging black hole video reconstruction, while baseline methods have difficulty accurately recovering temporal dynamics, our method demonstrates spatiotemporal structure that closely aligns with ground truth video (Fig. 5), under extremely sparse measurement conditions.
Figure 1: An overview of our proposed framework with Spatiotemporal Diffusion Priors, STeP, for scientific video inverse problems. Left: STeP combines the physic model of the target problem with a spatiotemporal diffusion prior that directly characterizes the video distribution. We show that such a prior can be efficiently obtained by fine-tuning a pre-trained image diffusion model with limited video data. Right: STeP can generate diverse solutions to a black hole video reconstruction problem that exhibit equally good fidelity with the measurements.
@@
Figure 2: A schematic comparison between prior works (top) and our STeP framework (bottom) for video inverse problems. The bold texts highlight the key differences between them. While prior works use an image diffusion model and enforce temporal consistency using simple heuristics or warping noise with optical flow, we directly learn a spatiotemporal diffusion prior.
@@ 2 Background
Video inverse problems
Existing approaches to video inverse problems have primarily focused on restoration and editing problems, such as super-resolution and JPEG artifact removal, for natural videos (chang2024how, ; daras2024warped, ; kwon2024visionxlhighdefinitionvideo, ; deng2025infiniteresolution, ; zou2025flair, ; kwon2025solving, ). Since it is computationally expensive to obtain a well-trained video diffusion model (VDM) for natural videos, existing methods instead rely on an image diffusion model (IDM) to process each video frame and propose various techniques to enforce temporal consistency. For example, the batch-consistent sampling (BCS) technique fixes the injected noise across all video frames (kwon2024visionxlhighdefinitionvideo, ; kwon2025solving, ). Another common approach, which we refer to as noise warping, first extracts optical flow from measurements and uses it to warp the injected noise across all video frames (chang2024how, ; daras2024warped, ; yeh2024diffir2vrzerozeroshotvideorestoration, ; deng2025infiniteresolution, ). As observed empirically, the dynamics of the injected noise translates into that of the generated video. These methods work well in restoration problems where the dynamics is mostly preserved in the measurements or video editing tasks where the dynamics is given. However, these methods face challenges when dealing with scientific problems because the measurements may belong to a different domain that is nontrivial to invert, or substantial spatiotemporal information may be lost in the measurement process.
Plug-and-play diffusion priors for inverse problems
Plug-and-play diffusion priors (PnPDP) constitute a family of methods that leverage diffusion models as priors for solving inverse problems. The goal is to draw samples from the posterior distribution $p({\mathbf{x}}_{0}|{\mathbf{y}})\,\propto\,p({\mathbf{y}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})$ where the likelihood $p({\mathbf{y}}|{\mathbf{x}}_{0})$ can be derived from knowledge of the physical system, and the prior $p({\mathbf{x}}_{0})$ is characterized by a diffusion model. Various formulations have been considered in the existing literature, including measurement guidance (chung2023diffusion, ; rout2023solving, ; kawar2022denoising, ; song2023pseudoinverseguided, ), variable splitting (coeurdoux2023plugandplay, ; li2024decoupleddataconsistencydiffusion, ; song2024solvinginverseproblemslatent, ; xu2024provably, ; wu2024principled, ; zhang2024improving, ), variational Bayes (mardani2023variational, ; feng2023score, ; feng2024variational, ), and Sequential Monte Carlo (SMC) (wu2023practical, ; trippe2023diffusion, ; cardoso2023monte, ; dou2024diffusion, ). A major advantage of PnPDP is its ability to handle inverse problems in a general way without task-specific design. Video inverse problems fit into the PnPDP framework in principle, but existing PnPDP methods have mainly focused on the static image setting. The main reason is that characterizing the prior distribution $p({\mathbf{x}}_{0})$ for videos is challenging due to their high dimensionality and a potentially limited number of samples for training. Prior work (kwon2024seeing, ) has explored the applicability of PnPDP to an optical scattering problem with a pixel-space video diffusion prior. In this work, we develop a general framework that can efficiently handle VIPs with larger spatiotemporal dimensions.
Diffusion models as image and video priors
Latent diffusion models (LDMs) (rombach2022high, ) offer a way to reduce the computational cost otherwise necessary to model the original high-dimensional data distribution. LDMs generate an efficient, low-dimensional latent representation ${\mathbf{z}}_{0}$ of data ${\mathbf{x}}_{0}$ with a pretrained perceptual compression encoder $\mathcal{E}$ and decoder $\mathcal{D}$, which satisfy ${\mathbf{z}}_{0}=\mathcal{E}({\mathbf{x}}_{0})$ and $\mathcal{D}({\mathbf{z}}_{0})\approx{\mathbf{x}}_{0}$. The compression models $\mathcal{E}$ and $\mathcal{D}$ can be trained as VAE variants (kingma2022autoencodingvariationalbayes, ; rezende2014stochasticbackpropagationapproximateinference, ; burgess2018understandingdisentanglingbetavae, ) with KL divergence regularization or VQGAN variants (gu2022vectorquantizeddiffusionmodel, ; esser2021tamingtransformershighresolutionimage, ; oord2018neuraldiscreterepresentationlearning, ) with quantization regularization. Video latent diffusion models are commonly believed to be hard to train due to the computational overhead of 3D modules and the requirement of a large video dataset (blattmann2023alignlatentshighresolutionvideo, ; zhou2023magicvideoefficientvideogeneration, ; zhang2024avidanylengthvideoinpainting, ; ho2022videodiffusionmodels, ; blattmann2023stablevideodiffusionscaling, ). Many recent methods tend to solve video modeling by fine-tuning from a pretrained IDM with a video dataset (wu2023tuneavideooneshottuningimage, ; blattmann2023alignlatentshighresolutionvideo, ; blattmann2023stablevideodiffusionscaling, ). Our work leverages these techniques to solve scientific VIPs.
@@ 3 Method
In this section, we introduce our framework, STeP, for solving video inverse problems with SpatioTemporal diffusion Priors. Figure 2 illustrates the conceptual difference between our approach and the prior works on video inverse problems. Designed mainly for restoration problems on natural videos, prior works rely on IDMs to incorporate spatial prior and enforce temporal consistency by injecting correlated noise in the diffusion process. As we will show in the experiments, these techniques do not work well when dealing with ill-posed scientific problems. Our method instead adopts a whole-video formulation (Sec. 3.1) and simultaneously handles spatial and temporal dimensions within a PnPDP framework (Sec. 3.2). The core of our method is to efficiently obtain spatiotemporal diffusion priors (Sec. 3.3), which directly learn temporal information from data and eliminate the need for designing task-specific temporal heuristics.
3.1 General approach
We consider the general formulation of video inverse problems where the goal is to recover a moving target ${\mathbf{x}}_{0}$ from the time-varying measurements
${\mathbf{y}}=\mathcal{A}({\mathbf{x}}_{0})+{\mathbf{n}}.$ (1)
Here $\mathcal{A}(\cdot)$ is the forward model, which could be nonlinear, and ${\mathbf{n}}$ is the measurement noise. In many scientific problems, such as black hole video reconstruction, substantial spatiotemporal information of ${\mathbf{x}}_{0}$ is lost in the measurement process due to the ill-posed and potentially nonlinear nature of $\mathcal{A}(\cdot)$, necessitating a prior on both the spatial and temporal dimensions of ${\mathbf{x}}_{0}$ for meaningful recovery.
Instead of taking a per-frame processing approach using IDMs, we propose to directly learn the video distribution $p({\mathbf{x}}_{0})$ from data and draw samples from the posterior distribution $p({\mathbf{x}}_{0}|{\mathbf{y}})\,\propto\,p({\mathbf{y}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})$ via the PnPDP framework. This enables us to deal with more general problems in which the extraction of temporal information directly from measurements is hard.
@@ 3.1 General approach
While the likelihood $p({\mathbf{y}}|{\mathbf{x}}_{0})$ can be derived from Equation (1), it is often challenging to characterize the prior distribution $p({\mathbf{x}}_{0})$ for videos due to its high dimensionality and potentially limited number of samples for training. To overcome these challenges, we propose to impose a spatiotemporal prior in the latent space. Assuming that the set of likely ${\mathbf{x}}_{0}$’s is in the range of a decoder $\mathcal{D}$, we have that $\exists\,{\mathbf{z}}_{0}$ s.t. ${\mathbf{x}}_{0}=\mathcal{D}({\mathbf{z}}_{0})$ and can thus rewrite Equation (1) as:
@@
It follows that the posterior $p({\mathbf{x}}_{0}|{\mathbf{y}})$ is the pushforward of the latent posterior $p({\mathbf{z}}_{0}|{\mathbf{y}})$ through $\mathcal{D}$, so it suffices to first generate latent samples from $p({\mathbf{z}}_{0}|{\mathbf{y}})$ and then decode them by $\mathcal{D}$.
3.2 Decoupled Annealing Posterior Sampling
STeP leverages the PnPDP framework to sample the posterior distribution given by Equation (2). In this work, we employ the Decoupled Annealing Posterior Sampling (DAPS) method (zhang2024improving, ), which has demonstrated strong performance on various scientific inverse problems zheng2025inversebench . It is also easily compatible with latent diffusion models, making it a suitable choice for our purpose.
DAPS samples the target latent posterior $p({\mathbf{z}}_{0}|{\mathbf{y}})$ by sequentially sampling $p({\mathbf{z}}_{t}|{\mathbf{y}})$ from $t=T$ to $t=0$. To do so, DAPS first generates a sample from $p({\mathbf{z}}_{T}|{\mathbf{y}})\approx\mathcal{N}({\bm{0}},\sigma_{\max}^{2}{\bm{I}})$ and sequentially draws a sample from $p({\mathbf{z}}_{t_{i-1}}|{\mathbf{y}})$ given a sample from $p({\mathbf{z}}_{t_{i}}|{\mathbf{y}})$ for $i=N,...,1$ based on a time schedule $\{t_{i}\}_{i=1}^{N}$. As shown by Proposition 1 of (zhang2024improving, ), this is possible if one can sample from
Since we have access to the gradient of this unnormalized distribution, we can sample it using MCMC methods, such as Langevin Monte Carlo (LMC) (welling2011bayesian, ) and Hamiltonian Monte Carlo (HMC) (betancourt2015hamiltonian, ). After obtaining ${\hat{\mathbf{z}}_{0}}\sim p({\mathbf{z}}_{0}|{\mathbf{z}}_{t_{i}},{\mathbf{y}})$, we then sample from $p({\mathbf{z}}_{t_{i-1}}|{\mathbf{y}})$ by sampling ${\mathbf{z}}_{t_{i-1}}\sim\mathcal{N}({\hat{\mathbf{z}}_{0}},\sigma_{t_{i-1}}^{2}{\bm{I}})$ due to Proposition 1 of (zhang2024improving, ). The pseudocode and more technical details of the STeP with DAPS are provided in Sec. A.
3.3 Efficient training of spatiotemporal diffusion priors
We propose to obtain spatiotemporal diffusion priors for scientific video inverse problems in an efficient way via the following three steps.